OpenStudy (nicole143):

Use summation nation to write each arithmetic series for the specified number of terms. Then evaluate the sum. 50 + 55 + 60 + ... ; n=7 What is summation nation? @mathmale @kirbykirby

OpenStudy (mathmale):

I see that each subsequent term of this arith. series is 5 greater than the preceding term. Agree with that?

OpenStudy (mathmale):

Since you need only 7 terms, you could obtain the remaining four in your head; the 4th is 65.

OpenStudy (nicole143):

Yes and okay, how do I write out the answer? Does it have to be in a certain format?

OpenStudy (nicole143):

Also, do they mean just in all or 7 after the 3 given?

OpenStudy (nicole143):

*mean 7 in all

OpenStudy (mathmale):

The first term is 50. We add 5 to that to get 55. The third term is 60, which is 55 + 5. Please try to predict the fourth thru 7th terms. All you need are the first 7 terms. Later the problem becomes a bit more complicated: we have to use summation notation to predict the nth sum.

OpenStudy (nicole143):

Okay, 50, 55, 60, 65, 70, 75, 80 And this one says to use summation nation.. Do I just write 50, 55, 60, 65, 70, 75, 80 on my page?

OpenStudy (mathmale):

Yes, although that would not be your final answer to this question. I thought it'd be helpful to write out these 7 terms so that you could add them up manually and thus know if the rest of your work on this problem is correct or not.

OpenStudy (nicole143):

So I don't need to write anything in summation nation?

OpenStudy (mathmale):

Here's what I would do, towards finding a formula for the nth term: 50+0(5) = 50 (first term) 50+1(5) = 55 (second term) 50+2(5) = 60 (third term) 50 +3(5) =65 (fourth term) If you see my pattern, continue my pattern to find the fifth term, please.

OpenStudy (nicole143):

50 + 4(5) = 70 50 + 5(5) = 75 50 + 6(5) = 80 50 + 7(5) = 85

OpenStudy (nicole143):

Opps just to 80

OpenStudy (mathmale):

Right. But that's OK. Recognizing the pattern is the key issue here. Nice work! Now let's change our findings into a formula for the nth term of this arith. sequence, OK?

OpenStudy (nicole143):

Okay, no idea how to do that ha

OpenStudy (mathmale):

the first term is 50. I wrote that as 50 + 0(5). Note that n = 1 signifies "first term." Note that that 0 can be written as 1-1. Thus, I propose that the formula for the first term is 50 + (1-1)(5). Does this produce the expected result, 50?

OpenStudy (mathmale):

in other words, Nicole, does 50 + (1-1)5 = 50 + 0(5)?

OpenStudy (nicole143):

Yes, it dose. So do I replace the 1-1 with "n" to signifies the replacement of the number for the equation?

OpenStudy (mathmale):

Exactly! would you do that, please? The formula for the nth term is ??

OpenStudy (nicole143):

50 + n(5) = An ?

OpenStudy (mathmale):

think carefully about what you are replacing with n. We don't replace all of that "1-1" with n. Test your theory. your formula says that the nth term is 50 + n(5). Let's see whether that predicts the first term properly. Let n = 1 in your equation. What is the result?

OpenStudy (nicole143):

oh, not 50.. Umm, 50 + n-1(5) = An ?

OpenStudy (mathmale):

If only you'll enclose that "n-1" inside parentheses, YES! Please do that. The formula for the nth term is ... ??

OpenStudy (nicole143):

50 + (n-1)5 = An

OpenStudy (nicole143):

Or 50 + (n-1)(5) = An ?

OpenStudy (nicole143):

Wait, no I take that ^^ back..

OpenStudy (mathmale):

Let's see what kirbykirby has to say!

OpenStudy (nicole143):

Okay :P

OpenStudy (mathmale):

In the meantime: What leads you to "take that back?"

OpenStudy (kirbykirby):

Hm well I just wanted to add that the question asks for a series, so shouldn't we be adding the terms, and thus use sigma notation, rather than just a sequence. Or maybe you are getting to that, which doesn't take much effort lol. But good job on the explanation so far mathmale

OpenStudy (nicole143):

FIOL I realized I would change the out come..

OpenStudy (mathmale):

I do believe we are 1) expressing a series in summation form, and 2) finding the actual sum of the series when n=7.

OpenStudy (nicole143):

I think so.. The very start of this post has everything I'm to do..

OpenStudy (mathmale):

Nicole, have you seen these formulas before?

OpenStudy (nicole143):

Parts look familiar but I didn't even know what summation nation was when we started this problem.. I don't think I could explain it now either

OpenStudy (mathmale):

\[\sum_{i-1}^{n}1=n;\sum_{i=1}^{n}i=\frac{ n(n+1) }{ 2 }??\] ... these are the formulas I was referring to.

OpenStudy (nicole143):

No.. I don't know

OpenStudy (mathmale):

The first formula says that if we add up n 1's, the sum is just n.

OpenStudy (nicole143):

Okay

OpenStudy (mathmale):

Will you have any contact with your teacher tomorrow? If so, i advise that you ask wehther or not you're expected to learn and know these particular summation formulas.

OpenStudy (nicole143):

I forget what the \[\sum_{ }^{ }\]Means..

OpenStudy (the_fizicx99):

I believe there's another formula

OpenStudy (nicole143):

I am, I just have a problem remembering and my teacher doesn't explain it in a way that I can easily pick up, where most of the time you guys and my tutor does..

OpenStudy (nicole143):

Oh, thats a bit confusing

OpenStudy (mathmale):

\[\sum_{i=1}^{n}1=n=1+1 + 1 + 1 + 1 + ....\] until you have added n 1's together.

OpenStudy (mathmale):

Yes, I made a mistake, which has now been corrected.

OpenStudy (mathmale):

\[\sum_{i=1}^{3}1=1+1+1=3=n\]is an easy example.

OpenStudy (nicole143):

Ohh, okay. I get it now.. I think ha

OpenStudy (mathmale):

I won't be able to provide a proof, but the following is also true:\[\sum_{i=1}^{n}i=\frac{ n(n+1) }{ 2 }\]

OpenStudy (mathmale):

Again, Nicole, as soon as you can, contact your teacher and find out if and when you need to know these "summation formulas". I would bet you do.

OpenStudy (nicole143):

So for my problem all that needs to be on the page is this right? 50 + 55 + 60 + 65 + 70 + 75 + 80 50 +0(5) = 50 ... 50 + 1(5) = 55 50 + (n-1)5 = An

OpenStudy (kirbykirby):

^Proof by induction :) haha but I think that would be too much lool

OpenStudy (nicole143):

Opps 50 +0(5) = 50 ... 50 + 6(5) = 80

OpenStudy (the_fizicx99):

I thought it was \[\sum_{k = 0}^{n-1} (a+kd) = \frac{ n }{ 2 }(2a(n-1)d)\]

OpenStudy (kirbykirby):

Oh I never remember such specific formulas :S Only the basic ones needed

OpenStudy (the_fizicx99):

typo, (2a + d(n-1))

OpenStudy (mathmale):

I believe there are different versions of the same general formula. I'm merely presenting the version with which I'm most familiar.

OpenStudy (nicole143):

Is that right?

OpenStudy (the_fizicx99):

Its really a matter of plugging in the values you already know, a, d and n

OpenStudy (mathmale):

I will let The Fizic answer that. He will first have to define what he means by a, d and n.

OpenStudy (nicole143):

So I need more for my page?

OpenStudy (mathmale):

Ladies and gentlemen, as much as I'd like to stick with this discussion, I need to hit the sack. Been on Open Study approx ten hours today.

OpenStudy (kirbykirby):

oh my @mathmale

OpenStudy (nicole143):

Okay, Thank you MM!

OpenStudy (mathmale):

So I bid you both a fond "good night." kirbykirby too!

OpenStudy (kirbykirby):

Good night!!

OpenStudy (the_fizicx99):

\[\sum_{k = 0}^{7-1}(50 + k * 5) = 7/2(2(50) + 5(7-1))\]

OpenStudy (nicole143):

@kirbykirby Is this all I need for my answer? 50 + 55 + 60 + 65 + 70 + 75 + 80 50 +0(5) = 50 ... 50 + 6(5) = 80 50 + (n-1)5 = An

OpenStudy (kirbykirby):

Nicole, yes a bit more is needed for your answer. Also, the question specifically asks for 7 terms.

OpenStudy (mathmale):

Buenas Noches, Luis! (Hey, that was an interesting expression; too bad you've erased it.

OpenStudy (kirbykirby):

Yes Fizi has the right formula down

OpenStudy (mathmale):

Over and out! Bye!

OpenStudy (nicole143):

There is 7, 50 to 80.

OpenStudy (the_fizicx99):

I did? goodnight Warren!!

OpenStudy (kirbykirby):

\[\sum_{i=0}^6 (50+5i)\]

OpenStudy (kirbykirby):

you can replace "i" with "k", it doesn't matter what letter is used.

OpenStudy (nicole143):

I am so confused.. I have 5 other problem I don't know how to do. So I think I'll just leave with one with what I have. If I only get partial point on the problem then so be it haha

OpenStudy (the_fizicx99):

xD

OpenStudy (kirbykirby):

do you know how to evaluate something like this: \[\sum_{i=1}^4 i\]

OpenStudy (nicole143):

I don't know what they E/Z thing is..

OpenStudy (nicole143):

*the

OpenStudy (kirbykirby):

it is a summation operator. It is basically a short form to write out a sum. So the sum I wrote above, it just saying "Add all the terms going from 1 to 4, by replacing "i" with 1, 2, 3, 4"

OpenStudy (kirbykirby):

Maybe some examples will help:

OpenStudy (kirbykirby):

\[\sum_{i=1}^4 i=1+2+3+4\]\[\sum_{i=1}^6 i=1+2+3+4+5+6\]\[\sum_{i=3}^7 i=3+4+5+6+7\]

OpenStudy (kirbykirby):

So more complicated expressions now: \[\sum_{i=1}^4 2i=2(1)+2(2)+2(3)+4(2)\]\[\sum_{i=1}^4 (5+i)=(5+1)+(5+2)+(5+3)+(5+4)\]

OpenStudy (kirbykirby):

Now, there is a formula, that mathmale mentioned: \[\sum_{i=1}^n i=\frac{n(n+1)}{2}\]. Why do we need this? Well, imagine if "n" was 100. We certainly would not want to add each number individually up to 100 (much too long to do so). So the formula above will gove you the sum of numbers from 1 to 100, if you replace "n" by 100.

OpenStudy (nicole143):

Okay, can you show me what mine would look like? @kirbykirby

OpenStudy (kirbykirby):

\[\sum_{i=0}^6 (50+5i)\] This is because it would give you these terms: [50+5(0)] + [50+50(1)] + [50+50(2)] + [50+50(3)] + [50+50(4)] + [50+50(5)] + [50+50(6)]

OpenStudy (kirbykirby):

even though it starts at 0 and ends at 6, please do realize that this gives you 7 terms. It might "look" unintuitive until you actually count the terms.

OpenStudy (nicole143):

Okay, thank you! I think I get it but I'm not sure. I have another problem with a different sequence, could you help me with that one as well?

OpenStudy (kirbykirby):

okay

OpenStudy (nicole143):

21 + 19 + 17 + ... ; n = 8 I know its by twos

OpenStudy (kirbykirby):

So, the first term... is 21 To get the 2nd term, using the first term, you do 21 - 2 Now, what do you do to get the 3rd term, 17

OpenStudy (nicole143):

21, 19, 17, 15, 13, 11, 9, 7

OpenStudy (kirbykirby):

Yes. But, we will want to write this in summation form. You want to construct ever number in your sequence based on the FIRST term. Like, for the previous question, we found every term of the sequence based on the first the 50, and see that we could get the n-th term by doing 50 + 5(n-1) (if n starts at 0)

OpenStudy (nicole143):

21 - 0(2) = 21 ?

OpenStudy (kirbykirby):

yes for the 1st term

OpenStudy (nicole143):

21 - 1(2) = 19 21 - 2(2) = 17 21 - 3(2) = 15 21 - 4(2) = 13 21 - 5(2) = 11 21 - 6(2) = 9 21 - 7(2) = 7